In the most recent year wholesome Foodsellers reported $6331 in sales with a 19% profit margin. The firmh as $ 15286 in assets. Next year the firm plans a 11% increase in sales. Assets and costs are proportional to sales. Debt and equity are not. The company doesnt pay dividends What is the external financing need for the next year? Select one: o 1549 o 3017 346 O 13951 2 479
Answer #1

Solution:

It is given that sales S0 = $6331 and growth = 11% so S1 = S0 *(1+11%) = 6331*(1+11%) = 7027.41

Asset = $15,286. Net profit margin = 19% and dividend payout = 0%

The formula for calculation of External fund needed is

EFN = Asset (Directly related to sales) / S0 * (S1-S0) - Liability (Directly related to sales) / S0 * (S1-S0) - Net profit margin * S1 * Retention rate

EFN = $15,286 / 6331 *(7027.41- 6331) - 0 (it is given that liability is not associated to sales) - 19% * 7027.41

EFN = 1681.46 - 1335.2079 = 346.25

So correct answer is option C) 346